# Ano : Faster is Better in Noisy Landscapes

### v1

@@ Abstract
Stochastic optimizers are central to deep learning, yet widely used methods such as Adam and Adan can degrade in non-stationary or noisy environments, partly due to their reliance on momentum-based magnitude estimates. We introduce Ano, a novel optimizer that decouples direction and magnitude: momentum is used for directional smoothing, while instantaneous gradient magnitudes determine step size. This design improves robustness to gradient noise while retaining the simplicity and efficiency of first-order methods. We further propose Anolog, which removes sensitivity to the momentum coefficient by expanding its window over time via a logarithmic schedule. We establish non-convex convergence guarantees with a convergence rate similar to other sign-based methods, and empirically show that Ano provides substantial gains in noisy and non-stationary regimes such as reinforcement learning, while remaining competitive on low-noise tasks such as standard computer vision benchmarks.
@@ 1 Introduction
Stochastic optimization is central to modern deep learning. Adaptive methods such as Adam (Kingma & Ba, 2015), and their variants (Reddi et al., 2018; Loshchilov & Hutter, 2019; Zaheer et al., 2018) are widely used because they automatically adjust step sizes and often accelerate early training. However, their behavior can degrade under noisy or non-stationary conditions: mini-batch stochasticity and data augmentation induce gradient noise (Mandt et al., 2017; Smith et al., 2018), labels may be ambiguous or noisy (Zhang et al., 2017; Song et al., 2022), and in reinforcement learning the training targets evolve over time (Henderson et al., 2019; Mnih et al., 2015). A common thread across these methods is their reliance on momentum estimates for both update direction and magnitude (Balles & Hennig, 2018). This entanglement makes them vulnerable: momentum magnitudes shrink in noisy regimes (Li et al., 2025), yielding overly conservative steps, or drift when noisy gradients accumulate (Bengio et al., 2021).
We propose A New Optimizer, abbreviated as Ano, which decouples direction and magnitude. Updates follow the sign of the momentum to benefit from its stability, while the step size is scaled by instantaneous gradient magnitudes, normalized by a variance term as in Adam. Intuitively, this design preserves the denoising benefits of momentum while enlarging the effective update to decorrelate the momentum.
@@
  • We evaluate Ano on supervised learning and deep reinforcement learning tasks, demonstrating substantial gains in noisy and non-stationary settings such as reinforcement learning, while remaining competitive in low-noise domains such as standard supervised benchmarks.
@@ 2 Related Work
AdaGrad (Duchi et al., 2011), AdaDelta (Zeiler, 2012) and RMSProp (Tieleman & Hinton, 2012) pioneered coordinate-wise adaptivity; Adam (Kingma & Ba, 2015) combined first- and second-moment estimates and became a default in deep learning. Yogi (Zaheer et al., 2018) stabilizes the second-moment accumulator for non-stationary regimes. More recently, Adan (Xie et al., 2024) couples adaptive moments with Nesterov-style lookahead and has emerged as a competitive baseline. Our optimizer Ano relates to this family through variance-aware step-size control, but differs in how direction and magnitude are constructed (see below).
@@
SignSGD and Signum (Bernstein et al., 2018) reduce updates to element-wise signs, offering scale invariance and communication efficiency in distributed settings. Lion (Chen et al., 2023) revisits sign-based updates with a tailored momentum schedule, yielding strong empirical results. Ano keeps the robustness of sign-informed directions but reintroduces gradient magnitudes through an explicit decoupling, trading pure scale invariance for finer adaptivity.
@@
Recent works such as Grams (Cao et al., 2025) decouple the update by using gradient signs for direction and the momentum norm for scaling. Ano adopts a complementary design: momentum provides a stable directional signal, while the raw gradient norm sets the step size. This hybridization aims to combine the resilience of sign-based directions with the adaptivity of moment estimators.
@@ Discussion.
Overall, Ano bridges sign-based and adaptive-moment families through a novel direction–magnitude decoupling strategy, where momentum is used to denoise and stabilize directions while gradient magnitudes preserve sensitivity to local curvature. This positioning explains its empirical robustness on high-variance RL tasks while remaining competitive in more stable regimes.
@@ 3 Algorithm
Motivation and Intuition.
In high-variance settings, optimizers like Adam rely on momentum to smooth updates through sharp minima and maxima. However, when gradient variance is high, the exponential moving average used for momentum can become dampened by noise, causing its magnitude to shrink and updates to become overly conservative. In addition, momentum may accumulate noisy, misaligned gradients, leading the update direction to drift away from the true descent path.
Momentum works best when it averages less correlated gradients that are at sufficiently different points in parameter space (Hermant et al., 2025). If updates are small, consecutive gradients are nearly identical and dominated by noise, making momentum unreliable. By using $|g_{k,i}|$ that tends to yield larger steps than $|m_{k,i}|$, we deliberately enlarge updates in noisy regions, while the normalization by $\sqrt{v_{k}}$ tempers outliers. This decorrelates successive gradients, enabling momentum to track the underlying descent direction more faithfully, while still permitting smooth progress in less noisy zones. Ano algorithm is presented below:
@@
 $v_{k}=v_{k-1}-(1-\beta_{2})\cdot\text{sign}(v_{k-1}-g_{k}^{2})\cdot g_{k}^{2}$
 $x_{k+1}=x_{k}-\frac{\eta_{k}}{\sqrt{v_{k}}+\epsilon}\cdot|g_{k}|\cdot\text{sign}(m_{k})-\eta_{k}\lambda x_{k}$
@@
Throughout this work, we employ Yogi’s additive update rule (Zaheer et al., 2018) to maintain the second-moment estimate:
$v_{k}=v_{k-1}-(1-\beta_{2})\,\mathrm{sign}(v_{k-1}-g_{k}^{2})\,g_{k}^{2},$
rather than Adam’s exponential moving average. This adjustment prevents abrupt increases in effective learning rates when $v_{t-1}$ significantly exceeds the current squared gradient, a known issue causing potential convergence instability. Empirically, this additive rule results in smoother loss curves, particularly in sparse-gradient scenarios, without additional memory or computational overhead. Thus, we adopt it as our default second-moment estimator.
Sign–Magnitude Decoupling.
Inspired by recent optimizers like Lion (Chen et al., 2023) and Signum (Bernstein et al., 2018), we explicitly decouple the direction and magnitude signals of parameter updates.
Concretely, recall that Adam updates parameters via
$x_{k+1}$ $=x_{k}-\frac{\eta}{\sqrt{v_{k}}+\epsilon}\cdot m_{k}=x_{k}-\frac{\eta}{\sqrt{v_{k}}+\epsilon}\,\underset{\text{magnitude}}{\underbrace{\bigl{|}m_{k}|}}\,\cdot\underset{\text{direction}}{\underbrace{\operatorname{sign}(m_{k})}}.$
$x_{k+1}=x_{k}-\frac{\eta}{\sqrt{v_{k}}+\epsilon}\,\underset{\text{magnitude}}{\underbrace{|g_{k}|}}\cdot\underset{\text{direction}}{\underbrace{\operatorname{sign}\!\bigl{(}m_{k})}}.$
@@
Since Ano relies solely on the momentum direction for updates, bias correction of its magnitude is unnecessary and omitted for simplicity (same as Lion). For the second moment estimator, we follow Yogi algorithm so we don’t use any bias correction. We therefore follow AdamW (Loshchilov & Hutter, 2019) and apply decoupled weight decay independently of gradient normalization.
@@
Like AdamW, Ano maintains first and second moments estimators : $m_{k}$ and $v_{k}$, each regulated by decay rates $\beta_{1}$ and $\beta_{2}$, with $\beta_{i}\in[0,1)$. We set $\beta_{1}=0.92$ and $\beta_{2}=0.99$ for stable convergence. Additionally, a weight decay coefficient $\lambda\in[0,+\infty)$ is employed to mitigate overfitting.
@@ 4 Extension
We define $\beta_{1,k}=1-\frac{1}{\log(k+2)}$, following theoretical and experimental motivations that favor slow and progressive adjustments to optimization hyperparameters. A slowly increasing $\beta_{1}$ progressively enlarges the effective averaging window of the momentum, reducing the influence of stochastic gradient noise as training progresses. In contrast, more aggressive schedules (e.g., square-root) can make the momentum insufficiently sensitive to recent gradient information, especially in the early iterations. We provide both empirical and ablation evidence in Section 7 to support the superiority of this scheduling choice against a square-root scheduling ($\beta_{1,k}=1-\frac{1}{\sqrt{k}}$) or a harmonic scheduling ($\beta_{1,k}=1-\frac{1}{k}$).
Full Anolog pseudo code can be found in Appendix A.
@@ 5.1 Theoretical Analysis
We provide non-asymptotic convergence guarantees for Ano under standard assumptions commonly used in adaptive stochastic optimisation (Kingma & Ba, 2015; Reddi et al., 2018). Consider the stochastic optimisation problem: $\min_{x\in\mathbb{R}^{d}}f(x),$ where $f$ is differentiable, $L$-smooth, and bounded below. Let $g_{k,i}$ denote the $i$-th coordinate of the stochastic gradient at iteration $k$ and $\mathcal{F}_{k-1}$ the filtration k-1. We assume that the gradient is bounded, $\|\nabla f(x)\|\leq G,\forall x\in\mathbb{R}^{d}$, the stochastic gradient is unbiased $(\mathbb{E}[g_{k,i}\mid\mathcal{F}_{k-1}]=\nabla_{i}f(x_{k}))$, and the variance is bounded $(\mathbb{E}[(g_{k,i}-\nabla_{i}f(x_{k}))^{2}\mid\mathcal{F}_{k-1}]\leq\sigma^{2})$.
@@
Following recent convergence analyses of sign-based optimizers, especially Lion (Dong et al., 2024), we assume a learning-rate schedule $\eta_{k}=\eta/k^{3/4}$ and $\beta_{1,k}=1-1/\sqrt{k}$, the iterates generated by Ano satisfy:
@@
$\mathbb{E}[f(x_{k+1})]\;\leq\;\mathbb{E}[f(x_{k})]\;-\frac{\eta_{k}}{\sqrt{2}G+\varepsilon}\,\mathbb{E}[\|\nabla f(x_{k})\|^{2}]\;+\mathcal{O}\Bigl{(}\frac{\eta_{k}}{k^{1/4}}\Bigr{)}\;+\mathcal{O}(\eta_{k}^{2}),$
@@ Discussion.
Our bound matches those recently established for sign-based optimizers such as Lion(Dong et al., 2024) and Signum(Bernstein et al., 2018), while relying on less restrictive assumptions (e.g., no requirement for growing batch sizes).Compared to adaptive schemes (SGD, Adam, Yogi) achieving $\mathcal{O}(K^{-1/2})$, our $\tilde{\mathcal{O}}(K^{-1/4})$ rate reflects a robustness–efficiency trade-off: Ano prioritizes stable sign-based updates, enforced by decaying schedules $\eta_{k}=\mathcal{O}(k^{-3/4})$ and $\beta_{1,k}=1-1/\sqrt{k}$. Full proofs are in Appendix C.
@@ 5.2 Noise Robustness Analysis
We assess noise robustness by training a CIFAR-10 CNN adding Gaussian noise $g_{k}\leftarrow g_{k}+\mathcal{N}(0,\sigma^{2})$ into every mini-batch gradient before the optimizer update. We vary only the noise level $\sigma$ over five values, keep each optimizer’s default $\beta$ and recommended learning rate (Full hyperparameters tabs can be found in Appendix B) for a computer vision task and report mean test accuracy over 5 seeds. [^1] Standard deviations omitted here for readability. The full table with standard deviation is available in Appendix D - tab 7
@@
Ano 82.25 78.78 71.03 65.96 59.81
Adam 80.89(-1.36) 76.08(-2.7) 66.43(-4.6) 60.56 (-5.4) 53.01(-6.8)
Lion 81.23(-1.02) 77.69(-1.09) 69.40(-1.63) 64.03(-1.93) 57.07(-2.74)
Grams 71.37(-10.88) 77.91(-0.87) 70.56(-0.47) 65.47(-0.49) 58.8(-1.01)
@@ 6.1 Computer Vision
Computer vision is a historically important domain for benchmarking optimization algorithms. In this section, we evaluate Ano on CIFAR-100 (Krizhevsky, 2009) using ResNet-34 (He et al., 2015).
@@
We use the standard CIFAR augmentation (random crop with 4-pixel padding + horizontal flip), following Zagoruyko & Komodakis (2017).
@@
Baseline
Adam $69.92\pm 0.18$ $0.042$
Adan $69.87\pm 0.09$ $0.052$
Lion $64.99\pm 2.22$ $0.100$
Grams $69.80\pm 0.10$ $0.047$
Ano $\mathbf{70.16}\pm\mathbf{0.047}$ $\mathbf{0.024}$
@@
Adam $69.85\pm 0.10$ $0.045$
Adan $70.53\pm 0.20$ $0.050$
Lion $69.61\pm 0.13$ $0.051$
Grams $69.82\pm 0.10$ $0.053$
Ano $\mathbf{71.15}\pm\mathbf{0.31}$ $\mathbf{0.021}$
@@
6.2 Transformers
As a cornerstone of modern artificial intelligence, natural language processing (NLP) warrants careful evaluation of optimization algorithms. The unique challenges of NLP, such as large parameter spaces, sparse gradients, and label noise, can affect optimizer performance (Mosbach et al., 2021). We assess the effectiveness of Ano by comparing it to standard baselines on the GLUE benchmark (Wang et al., 2019), covering eight sentence- and sentence-pair classification tasks (excluding WNLI, following standard practice due to its unreliable performance).

### v2

@@ Abstract
Stochastic optimizers are central to deep learning, yet widely used methods such as Adam and Adan can degrade in non-stationary or noisy environments, partly due to their reliance on momentum-based magnitude estimates. We introduce Ano, a novel optimizer that decouples direction and magnitude: momentum is used for directional smoothing, while instantaneous gradient magnitudes determine step size. This design improves robustness to gradient noise while retaining the simplicity and efficiency of first-order methods. We further propose Anolog, which removes sensitivity to the momentum coefficient by expanding its window over time via a logarithmic schedule. We establish non-convex convergence guarantees with a convergence rate similar to other sign-based methods, and empirically show that Ano provides substantial gains in noisy and non-stationary regimes such as reinforcement learning, while remaining competitive on low-noise tasks.
@@ 1 Introduction
Stochastic optimization is central to modern deep learning. Adaptive methods such as Adam (kingma2015adam), and their variants (reddi2018convergence; loshchilov2018decoupled; Zaheer2018Adaptive) are widely used because they automatically adjust step sizes and often accelerate early training. However, their behavior can degrade under noisy or non-stationary conditions: mini-batch stochasticity and data augmentation induce gradient noise (mandt2017stochastic), labels may be ambiguous or noisy (zhang2017understanding; song2022learning), and in reinforcement learning the training targets evolve over time (henderson2018matters; mnih2016asynchronous). A key limitation is that Adam couples update direction and magnitude through momentum: prior work (balles2018dissecting) shows that the momentum sign already captures most directional information, while its magnitude and the second-moment estimate impose heavy smoothing.
We propose A New Optimizer, abbreviated as Ano, designed to handle noisy optimization landscapes. First, we decouple direction and magnitude: updates follow the momentum sign for stability, while the step size is scaled by an SNR-like ratio of instantaneous gradients, avoiding the sluggishness of momentum-based magnitudes. Second, we revisit Yogi’s asymmetric variance update (Zaheer2018Adaptive), which accelerates recovery after noise spikes, and introduce an additional decay factor to control its memory. This preserves Yogi’s fast responsiveness while ensuring smoother adaptation under highly stochastic gradients.
@@
  • We evaluate Ano on supervised and reinforcement learning tasks, showing clear gains in noisy, non-stationary settings while remaining competitive in standard benchmarks.
@@ 2 Related Work
AdaGrad (duchi2011adaptive), AdaDelta (zeiler2012adadeltaadaptivelearningrate) and RMSProp (tieleman2012rmsprop) pioneered coordinate-wise adaptivity; Adam (kingma2015adam) combined first- and second-moment estimates and became a default in deep learning. Yogi (Zaheer2018Adaptive) stabilizes the second-moment accumulator for non-stationary regimes. More recently, Adan (xie2024adan) couples adaptive moments with Nesterov-style lookahead and has emerged as a competitive baseline. Our optimizer Ano relates to this family through variance-aware step-size control, but differs in how direction and magnitude are constructed.
@@
SignSGD and Signum (bernstein2018signsgd) reduce updates to element-wise signs, offering scale invariance and communication efficiency in distributed settings. Lion (chen2023symbolic) revisits sign-based updates with a tailored momentum schedule, yielding strong empirical results. Ano keeps the robustness of sign-informed directions but reintroduces gradient magnitudes through an explicit decoupling, trading pure scale invariance for finer adaptivity.
@@
Recent works such as Grams (cao2024grams) decouple the update by using gradient signs for direction and the momentum norm for scaling. Ano adopts a complementary design: momentum provides a stable directional signal, while the raw gradient norm sets the step size. This hybridization aims to combine the resilience of sign-based directions with the adaptivity of moment estimators.
Optimization under non-stationarity.
Non-stationarity is a known stressor for optimization, particularly in RL. Prior work tackled it with task-specific procedures such as Normalize and Project (NaP) (lyle2024normalization) and meta-learned optimizers for RL (lan2025learning). While these highlight the need for robustness to evolving objectives, they are not first-order per-parameter adaptive optimizers in the Adam/Lion sense. Ano instead offers a simple, general optimizer that retains such efficiency while improving stability under noise and non-stationarity.
@@ Discussion.
Ano unifies sign-based and adaptive-moment ideas via a per-parameter direction–magnitude split (momentum for direction, raw gradients for scale), which we find particularly robust in high-variance regimes while remaining competitive on standard tasks.
@@ 3 Algorithm
The full Ano algorithm is summarized in Algorithm 1. Like Adam, it maintains first- and second-moment estimates $m_{k}$, $v_{k}$, but introduces two key innovations described below: one targeting the decoupling of update direction and magnitude, and the other improving variance adaptation under noisy gradients.
Ano algorithm is presented below:
@@
 $v_{k}=\beta_{2}v_{k-1}-(1-\beta_{2})\cdot\text{sign}(v_{k-1}-g_{k}^{2})\cdot g_{k}^{2}$
 $\hat{v_{k}}=\frac{v_{k}}{1-\beta 2^{k}}$
 $x_{k+1}=x_{k}-\frac{\eta_{k}}{\sqrt{\hat{v_{k}}}+\epsilon}\cdot|g_{k}|\cdot\text{sign}(m_{k})-\eta_{k}\lambda x_{k}$
@@
Sign–Magnitude Decoupling.
We explicitly decouple the direction and magnitude of parameter updates to mitigate the conservative dynamics of Adam. In Adam, both signals are derived from the momentum term $m_{k}$, so when large noise spikes occur, their opposing effects can partially cancel out, reducing the effective momentum and thereby slowing down the updates. Ano keeps the direction $\operatorname{sign}(m_{k})$ for robustness to noise but replaces the momentum magnitude with the instantaneous gradient norm $|g_{k}|$ for a better scaling.
Concretely, recall that Adam updates parameters via
$x_{k+1}$ $=x_{k}-\frac{\eta_{k}}{\sqrt{v_{k}}+\epsilon}\cdot m_{k}=x_{k}-\frac{\eta_{k}}{\sqrt{\hat{v_{k}}}+\epsilon}\,\underset{\text{magnitude}}{\underbrace{\bigl|m_{k}|}}\,\cdot\underset{\text{direction}}{\underbrace{\operatorname{sign}(m_{k})}}.$
$x_{k+1}=x_{k}-\frac{\eta_{k}}{\sqrt{\hat{v_{k}}}+\epsilon}\,\underset{\text{magnitude}}{\underbrace{|g_{k}|}}\cdot\underset{\text{direction}}{\underbrace{\operatorname{sign}\!\bigl(m_{k})}}.$
@@
Ano improves variance dynamics for stability and fast recovery in particular for non-stationnary landscape. Adam’s exponential moving average (kingma2015adam) keeps noise spikes alive for many iterations, inflating the variance estimate and shrinking steps even after the signal improves. Yogi (Zaheer2018Adaptive) addresses this with asymmetric updates for faster decay. We extend Yogi by introducing a decay factor that explicitly controls variance memory, maintaining the exponential structure while allowing smooth forgetting of outdated information. This mechanism naturally assigns greater weight to recent gradients, thereby enhancing adaptation in dynamic environments. which is essential in non-stationary environments. Formally,
$v_{k}=\beta_{2}v_{k-1}-(1-\beta_{2})\,\mathrm{sign}(v_{k-1}-g_{k}^{2})\,g_{k}^{2},$
turning the variance term into a memory-controlled statistic rather than a purely reactive estimate.
@@
Since Ano relies solely on the momentum direction for updates, bias correction of its magnitude is unnecessary and omitted for simplicity (same as Lion) but keep it for the variance estimate. Weight decay follows AdamW (loshchilov2018decoupled) for decoupled regularization.
@@
Like Adam, Ano maintains first and second moments estimators : $m_{k}$ and $v_{k}$, each regulated by decay rates $\beta_{1}$ and $\beta_{2}$, with $\beta_{1}\in[0,1)$ and $\beta_{2}\in[\frac{1}{2},1)$. We set $\beta_{1}=0.92$ and $\beta_{2}=0.99$ for stable convergence. Additionally, a weight decay coefficient $\lambda\in[0,+\infty)$ is employed to mitigate overfitting.
@@ 4 Extension
We define $\beta_{1,k}=1-\frac{1}{\log(k+2)}$, motivated by both theoretical considerations and empirical evidence favoring slow, progressive adjustments to optimization hyperparameters. A gradually increasing $\beta_{1}$ enlarges the effective averaging window of the momentum, thereby reducing the impact of stochastic gradient noise as training proceeds. In contrast, more aggressive schedules (e.g., square-root) may render the momentum insufficiently responsive to recent gradient information, particularly in non-stationary settings where rapid adaptation is crucial. Section 7 provides empirical and ablation results comparing this logarithmic schedule against square-root ($\beta_{1,k}=1-\frac{1}{\sqrt{k}}$) and harmonic ($\beta_{1,k}=1-\frac{1}{k}$) schedules.
Full Anolog pseudo code can be found in Appendix A - Algorithm 2.
@@ 5.1 Theoretical Analysis
We provide non-asymptotic convergence guarantees for Ano under standard assumptions commonly used in adaptive stochastic optimisation (kingma2015adam; reddi2018convergence). Consider the stochastic optimisation problem: $\min_{x\in\mathbb{R}^{d}}f(x),$ where $f$ is differentiable, $L$-smooth, and bounded below. Let $g_{k,i}$ denote the $i$-th coordinate of the stochastic gradient at iteration $k$ and $\mathcal{F}_{k-1}$ the filtration k-1. We assume that the gradient is bounded, $|\nabla_{i}f(x_{k})|\leq G,\forall x\in\mathbb{R}^{d}$, the stochastic gradient is unbiased $(\mathbb{E}[g_{k,i}\mid\mathcal{F}_{k-1}]=\nabla_{i}f(x_{k}))$, and the variance is bounded $(\mathbb{E}[(g_{k,i}-\nabla_{i}f(x_{k}))^{2}\mid\mathcal{F}_{k-1}]\leq\sigma^{2})$.
@@
Following recent convergence analyses of sign-based optimizers, especially Lion (dong2024convergence), we assume a learning-rate schedule $\eta_{k}=\eta/k^{3/4}$ and $\beta_{1,k}=1-1/\sqrt{k}$, the iterates generated by Ano satisfy:
@@
$\mathbb{E}[f(x_{k+1})]\;\leq\;\mathbb{E}[f(x_{k})]\;-\frac{\eta_{k}}{\tilde{G}+\varepsilon}\,\mathbb{E}[\|\nabla f(x_{k})\|^{2}]\;+\mathcal{O}\Bigl(\frac{\eta_{k}}{k^{1/4}}\Bigr)\;+\mathcal{O}(\eta_{k}^{2}),$
@@ Discussion.
Our bound matches those recently established for sign-based optimizers such as Lion(dong2024convergence) and Signum(bernstein2018signsgd), while relying on less restrictive assumptions (e.g., no requirement for growing batch sizes). Compared to adaptive schemes (SGD, Adam, Yogi) achieving $\mathcal{O}(K^{-1/2})$, our $\tilde{\mathcal{O}}(K^{-1/4})$ rate stems from a fundamental limitation of sign-based methods: ensuring stable updates requires decaying step sizes $\eta_{k}=\mathcal{O}(k^{-3/4})$ which, in turn, constrains the overall convergence rate. Full proofs are in Appendix C.
@@ 5.2 Noise Robustness Analysis
We assess noise robustness by training a CIFAR-10 CNN adding Gaussian noise $g_{k}\leftarrow g_{k}+\mathcal{N}(0,\sigma^{2})$ into every mini-batch gradient before the optimizer update (krizhevsky2009learning). We vary only the noise level $\sigma$ over five values, keep each optimizer’s default $\beta$ and recommended learning rate (Full hyperparameters tabs can be found in Appendix B) for a computer vision task and report mean test accuracy over 5 seeds.11195%CI omitted here for readability. The full table with 95% CI is available in Appendix D - tab 7
@@
Ano 82.10 78.71 70.88 65.93 59.54
Adam 80.67(-1.43) 75.97(-2.74) 66.86(-4.02) 60.83(-5.10) 52.46(-7.08)
Lion 81.04(-1.05) 77.80(-0.91) 69.62(-1.26) 64.02(-1.91) 56.82(-2.72)
Grams 71.34(-10.76) 77.90(-0.81) 70.57(-0.31) 65.47(-0.46) 58.80(-0.74)
@@ 6.1 Computer Vision
Computer vision is a historically important domain for benchmarking optimization algorithms. In this section, we evaluate Ano on CIFAR-100 (krizhevsky2009learning) using ResNet-34 (he2016deep).
@@
We use the standard CIFAR augmentation (random crop with 4-pixel padding + horizontal flip), following zagoruyko2017wideresidualnetworks.
@@
Default
Adam $69.57\pm 0.22$ $0.037$
Adan $69.87\pm 0.09$ $0.049$
Lion $66.25\pm 0.61$ $0.064$
Grams $68.33\pm 0.40$ $0.045$
Ano $\mathbf{70.31}\pm\mathbf{0.50}$ $\mathbf{0.015}$
Anolog $64.84\pm 1.19$ $0.019$
@@
Adam $69.61\pm 0.23$ $0.042$
Adan $69.09\pm 0.16$ $0.049$
Lion $68.77\pm 0.26$ $0.048$
Grams $68.11\pm 0.27$ $0.048$
Ano $\mathbf{69.89}\pm\mathbf{0.42}$ $\mathbf{0.022}$
Anolog $68.41\pm 0.58$ $\mathbf{0.032}$
@@
6.2 Natural Language Processing
As a cornerstone of modern artificial intelligence, natural language processing (NLP) warrants careful evaluation of optimization algorithms. The unique challenges of NLP, such as large parameter spaces, sparse gradients, and label noise, can affect optimizer performance (mosbach2021on). We assess the effectiveness of Ano by comparing it to standard baselines on the GLUE benchmark (wang2019gluemultitaskbenchmarkanalysis), covering eight sentence- and sentence-pair classification tasks (excluding WNLI, following standard practice due to its unreliable performance).